Transform the following equation by completing the square. Identify the conic section represented by each equation or its boundary.
x=0.2y^2-2y-3

Answer by stanbon(75887)

Transform the following equation by completing the square. Identify the conic section represented by each equation or its boundary.
x=0.2y^2-2y-3
0.2y^2-2y = x+3
--------------
0.2[y^2 - 10y + ?] = x + 3 + 0.2*?
---
0.2[y^2-10y+25] = x + 3 + 0.2*25
0.2(y-5)^2 = x + 3 + 5
---
x = 0.2(y-5)^2 - 8
----
Parabola opening to the right.
=================================
Cheers,
Stan H.
==================